# Asymptotic confidence bands for the histogram regression estimator

### v1

@@ 4.3 Moment bounds for binomial distributions
If (4) is satisfied, we have
$\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{p}_{\delta}(x)}{p_{\delta}(x)}-1\Big{|}=o_{\mathbb{P}}\big{(}(\log n)^{-3/2})\big{)}.$
@@
$\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{p}_{\delta}(x)}{p_{\delta}(x)}-1\Big{|}$ $=$ $\sup_{x\in[0,1]^{p}}\frac{1}{np_{\delta}(x)}\Big{|}\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}-np_{\delta}(x)\Big{|}.$
Since $x\mapsto\hat{p}_{\delta}(x)$ is piecewise constant with at most $\delta{-p}$ different values, the supremum can be handled with a union bound. Moreover, we can exploit that $\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}$ is $Bin(n,p_{\delta}(x))$-distributed. The binomial moment result (26) and (7) yield for any $q\in\mathbb{N}$ that
$\mathbb{P}\left(\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{p}_{\delta}(x)}{p_{\delta}(x)}-1\Big{|}\geq\frac{\kappa}{(\log n)^{3/2}}\right)$
$\qquad\leq\delta^{-p}\frac{1}{(nc_{X}\delta^{p})^{q}}\mathbb{E}\Big{[}\Big{|}\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\delta}(x)\}-np_{\delta}(x)\Big{|}^{q}\Big{]}\frac{(\log n)^{3q/2}}{\kappa^{q}}$
$\qquad\lesssim\frac{\delta^{-p}(\log n)^{3q/2}}{\kappa^{q}(\delta^{p}n)^{q/2}}=o(1)$
@@
Appendix A Approximation of suprema of empirical processes
@@
Suppose also that for some $b\geq\tilde{\sigma}>0,$ and $\nu\in[4,\infty]$, we have $\sup_{f\in\mathcal{F}}P|f|^{k}\leq\tilde{\sigma}^{2}b^{k-2}$ for $k=2,3$ and $\|F\|_{P,\nu}\leq b$. Then for every $\gamma\in(0,1)$ there exists a random variable $\mathbf{S}\overset{d}{=}\sup_{f\in\mathcal{F}}G_{P}f$ such that
$\mathbb{P}\Big{(}\Big{|}\sup_{f\in\mathcal{F}}\mathbb{G}_{n}f-\mathbf{S}\Big{|}>\frac{b\Gamma_{n}}{\gamma^{1/2}n^{1/2-1/\nu}}+\frac{(b\tilde{\sigma})^{1/2}\Gamma_{n}^{3/4}}{\gamma^{1/2}n^{1/4}}+\frac{(b\tilde{\sigma}^{2}\Gamma_{n}^{2})^{1/3}}{\gamma^{1/3}n^{1/6}}\Big{)}\leq C\Big{(}\gamma+\frac{\log n}{n}\Big{)}$
where $)\Gamma_{n}=cv(\log n\vee\log(Ab/\tilde{\sigma}$)), and $c,C>0$ are constants that depend only on $\nu$ ($1/\nu$ is interpreted as 0 when $\nu=\infty$).

### v2

@@ 6.3 Moment bounds for binomial distributions
Let Assumption 2.1 and (6) be satisfied. Then we have
$\sup_{x\in\mathcal{X}}\Big|\frac{\widehat{p}_{\Delta}(x)}{p_{\Delta}(x)}-1\Big|=o_{\mathbb{P}}\big((\log n)^{-3/2})\big).$
@@
$\sup_{x\in\mathcal{X}}\Big|\frac{\widehat{p}_{\Delta}(x)}{p_{\Delta}(x)}-1\Big|$ $=\sup_{x\in\mathcal{X}}\frac{1}{np_{\Delta}(x)}\Big|\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}-np_{\Delta}(x)\Big|.$
Since $x\mapsto\widehat{p}_{\Delta}(x)$ is piecewise constant with at most $\Delta$ different values, the supremum can be handled with a union bound. Moreover, we can exploit that $\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}$ is $Bin(n,p_{\Delta}(x))$-distributed. The binomial moment result (33) and (10) yield for any $q\in\mathbb{N}$ that
$\mathbb{P}\left(\sup_{x\in\mathcal{X}}\Big|\frac{\widehat{p}_{\Delta}(x)}{p_{\Delta}(x)}-1\Big|\geq\frac{\kappa}{(\log n)^{3/2}}\right)$
$\qquad\leq\frac{\Delta}{(nc_{X}c_{p}/\Delta)^{q}}\mathbb{E}\Big[\Big|\sum_{i=1}^{n}\mathbb{I}\{X_{i}\in A_{\Delta}(x)\}-np_{\Delta}(x)\Big|^{q}\Big]\frac{(\log n)^{3q/2}}{\kappa^{q}}$
$\qquad\lesssim\frac{\Delta(\log n)^{3q/2}}{\kappa^{q}(n/\Delta)^{q/2}}=o(1)$
@@
Appendix A Extensions
For the general partition in Assumption 2.1 we chose disjoint cells, but intersecting cells are also possible. For this we fix a mapping $x\mapsto A_{\Delta}(x)$, where all $(A_{\Delta}(x))_{x\in\mathcal{X}}$ satisfy (3), and we assume that $\{A_{\Delta}(x)\mid x\in\mathcal{X}\}$ is a subset of some VC-class of sets, not depending on $\Delta$. For instance the cells $A_{\Delta}(x)$ can be chosen as balls (or hypercubes) with middle point $x$ and radius (side length) depending on $\Delta$. We use the same notation for the estimator $\widehat{m}$ in (4) and the assumptions in Theorem 2.2. However, $\Delta$ is not the number of cells anymore, and we do not assume that there are only finitely many cells for each $\Delta$. Thus the parts of the proof applying the union bound and the finite cardinality of the function class $\mathcal{F}_{\Delta}$ have to be modified. The general proof structure is still the same as in Sections 3 and 6.
One starts with the same expansion with terms $T_{1}$, $T_{2}$ and $T_{3}$ from (21), (22) and (23), but with an adjusted $\mathbf{S}_{\Delta}$ discussed below. To handle $T_{1}$, (25) works out the same way, but for (24) and showing negligibility of $T_{3}$ a different argument is needed. Let $P$ denote the distribution of $X_{1}$ and $P_{n}$ the empirical distribution based on $X_{1},\dots,X_{n}$. Then one can write
$\sup_{x\in\mathcal{X}}|\widehat{p}_{\Delta}(x)-p_{\Delta}(x)|=\sup_{g\in\mathcal{G}_{n}}|P_{n}g-Pg|$
for the function class $\mathcal{G}_{n}=\{\mathbb{I}_{A_{\Delta}(x)}\mid x\in\mathcal{X}\}$ (with $\Delta=\Delta_{n}$), which is a subset of a VC-class not depending on $n$. We have $|g|\leq 1$ and $(Pg^{2})^{1/2}=(p_{\Delta}(x))^{1/2}\leq(C_{X}C_{p})^{1/2}\Delta^{-1/2}$ for all $g\in\mathcal{G}_{n}$ by (10) and (3). Applying Theorem 37 in Pollard, (1984, p. 34), one obtains
$\sup_{x\in\mathcal{X}}|\widehat{p}_{\Delta}(x)-p_{\Delta}(x)|=o_{\mathbb{P}}(\alpha_{n}\Delta^{-1})$
for each positive non-increasing sequence $\alpha_{n}$ with $(\Delta\log n)/(n\alpha_{n}^{2})=o(1)$. In view of (6) we can choose $\alpha_{n}=(\log n)^{-2}$. With this one also obtains a similar result as (27) because
$\sup_{x\in\mathcal{X}}\left|p_{\Delta}(x)\left(\frac{1}{\widehat{p}_{\Delta}(x)}-\frac{1}{p_{\Delta}(x)}\right)\right|\mathbb{I}\{\widehat{p}_{\Delta}(x)>0\}$
$\leq$ $\frac{\sup_{x\in\mathcal{X}}|\widehat{p}_{\Delta}(x)-p_{\Delta}(x)|}{\inf_{x\in\mathcal{X}}p_{\Delta}(x)}\sup_{x\in\mathcal{X}}\frac{p_{\Delta}(x)}{p_{\Delta}(x)-\sup_{t\in\mathcal{X}}|\widehat{p}_{\Delta}(t)-p_{\Delta}(t)|}=o_{\mathbb{P}}((\log n)^{-2}).$
Further for the term $\mathbb{I}\{\widehat{p}_{\Delta}(x)=0\}$ in the upper bound of $T_{3}$ note that for $\eta>0$
$\mathbb{P}\left((\log n)^{3/2}\sup_{x\in\mathcal{X}}\mathbb{I}\{\widehat{p}_{\Delta}(x)=0\}>\eta\right)\leq$ $\mathbb{P}\left(\exists x\in\mathcal{X}:\widehat{p}_{\Delta}(x)=0\right)$
$\leq$ $\mathbb{P}\left(\sup_{x\in\mathcal{X}}|\widehat{p}_{\Delta}(x)-p_{\Delta}(x)|\geq c_{X}c_{p}\Delta^{-1}\right)=o(1),$
and to handle term (24) for $T_{1}$ works in the same way. Thus $T_{1}$ is negligible. For negligibility of $T_{3}$ one also needs the rate $\mathbf{S}_{\Delta}=\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})$, and arguments for this are based on the cell structures. We consider this in the examples below. The dominating term $T_{2}$ can be treated as in Theorem 3.1. For this let $\mathcal{A}_{\Delta}(x)$ be the VC-class containing all sets $A_{\Delta}(x)$. Then $\mathcal{F}_{\Delta}$ with the analogous definition as in (19) and envelope $F$ from (30) is a subset of the VC-type class
$\{(t,\epsilon)\mapsto|\epsilon|c\Delta^{1/2}\mathbb{I}_{A}\mid A\in\mathcal{A}_{\Delta}(x),c\in[0,(c_{X}c_{\sigma^{2}}c_{p})^{-1/2}],x\in\mathcal{X}\}$
such that the covering number is bounded by
$\sup_{Q\in\mathcal{Q}}N(\mathcal{F}_{\Delta},\|\cdot\|_{Q,2},\kappa\|F\|_{Q,2})\leq C\kappa^{-V}$
for some $C$ and $V$ not depending on $\Delta$. So we even get a better upper bound than in (31) and the remaining part of the proof of Theorem 3.1 can be done analogously. The approximation has the structure
$\mathbf{S}_{\Delta}\overset{d}{=}\sup_{x\in\mathcal{X}}|B_{\Delta}(f_{x,\Delta})|,$
where the centered Gaussian process $B_{\Delta}$ has the covariance formula
$\mathrm{Cov}(B_{\Delta}(f_{x_{1},\Delta}),B_{\Delta}(f_{x_{2},\Delta}))=\frac{\int_{A_{\Delta}(x_{1})\cap A_{\Delta}(x_{2})}\sigma^{2}(t)f_{X}(t)\,dt}{(s_{\Delta}(x_{1})s_{\Delta}(x_{2}))^{1/2}},$
and the same distribution as the process
$\frac{1}{(s_{\Delta}(x))^{1/2}}\int\sigma(t)(f_{X}(t))^{1/2}\mathbb{I}_{A_{\Delta}(x)}(t)\,dW(t)$ (35)
with a $p$-parameter Brownian sheet $W$. Due to the overlapping cells, we don’t have an diagonal covariance structure anymore and $\mathbf{S}_{\Delta}$ cannot be represented by a finite number of Gaussian random variables anymore.
Example A.1.
Consider the homoscedastic case with $X\sim U[0,1]^{p}$. Then the covariance structure simplifies to
$\mathrm{Cov}(B_{\Delta}(f_{x_{1},\Delta}),B_{\Delta}(f_{x_{2},\Delta}))=\frac{\operatorname{vol}(A_{\Delta}(x_{1})\cap A_{\Delta}(x_{2}))}{(\operatorname{vol}(A_{\Delta}(x_{1}))\operatorname{vol}(A_{\Delta}(x_{2})))^{1/2}},$
and quantiles of $\mathbf{S}_{\Delta}$ can be directly simulated to obtain a confidence band. For the proof discussed above one needs the convergence rate of $\mathbf{S}_{\Delta}$. Let us consider the case where $A_{\Delta}(x)$ is a ball around $x$ (intersected with $0,1]^{p}$) with radius $r_{\Delta}(x)$, which may depend on $x$ with the assumption
$c_{r}\Delta^{-1/p}\leq r_{\Delta}(\cdot)\leq C_{r}\Delta^{-1/p}$
for some constants $C_{r}\geq c_{r}>0$. Define the distance between $x_{1},x_{2}\in[0,1]^{p}$ as
$d(x_{1},x_{2})=\big(\mathrm{Var}(B_{\Delta}(f_{x_{1},\Delta})-B_{\Delta}(f_{x_{2},\Delta}))\big)^{1/2}=\left(1-\frac{\operatorname{vol}(A_{\Delta}(x_{1})\cap A_{\Delta}(x_{2}))}{(\operatorname{vol}(A_{\Delta}(x_{1}))\operatorname{vol}(A_{\Delta}(x_{2})))^{1/2}}\right)^{1/2},$
and assume it can be upper bounded by
$d(x_{1},x_{2})\leq\left(c\|x_{1}-x_{2}\|\Delta^{1/p}\right)^{1/2}$
for some constant $c$. For instance for dimension $p=3$ one can apply a sphere-sphere intersection formula and assume Lipschitz continuity of the map $x\mapsto r_{\Delta}(x)$. Then one obtains an upper bound for the covering number $N(\epsilon,\mathcal{F}_{\Delta},d)\lesssim(c\Delta^{1/p}/\epsilon^{2})^{p}$. By Massart, (2007, Theorem 3.18) we obtain
$\mathbb{E}[\mathbf{S}_{\Delta}]\lesssim\int_{0}^{1}\sqrt{\log((c\Delta^{1/p}/\epsilon^{2})^{p})}\,d\epsilon=\mathcal{O}(\sqrt{\log\Delta})$
and thus $\mathbf{S}_{\Delta}=\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})$.
Example A.2.
Consider the special case where $t\in A_{\Delta}(x)$ is equivalent to $(x-t)/\delta\in A$ for one fixed set $A$ such that one can write $\mathbb{I}\{t\in A_{\Delta}(x)\}=K(\frac{x-t}{\delta})$ with $K=\mathbb{I}_{A}$ and for example $A=[-1,1]^{p}$. For the assumptions on $\Delta$ as in Example 2.3 let $\Delta=\delta^{-p}$. The estimator $\widehat{m}$ in (4) is then a Nadaraya-Watson estimator with kernel $K$ and bandwidth $\delta$. As in the last steps of classical confidence band proof structures, see Chao et al., (2017), one can now replace $\sigma(t)(f_{X}(t))^{1/2}$ inside the integral (35) by $\sigma(x)(f_{X}(x))^{1/2}$ and use a factor function to obtain the simple process
$Z_{\delta}(x)=\frac{1}{\sqrt{\delta^{p}}}\int K\left(\frac{x-t}{\delta}\right)\,dW(t).$
For the replacement some assumptions on the functions $\sigma$ and $f_{X}$ are needed. For the proof structure considered above note that instead of the rate of $\mathbf{S}_{\Delta}$ one can use the rate $\sup_{x}|Z_{\delta}(x)|=\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})$, and one obtains the limit distribution with the classical Rosenblatt, (1976) result. We are not aware of literature about asymptotic confidence bands based on Nadaraya-Watson estimators in the multidimensional covariate case for regression functions with Hölder regularity $\alpha\in(0,1]$. This additionally demonstrates the usefulness of the new proof structure.
Appendix B Approximation of suprema of empirical processes
@@
Suppose also that for some $b\geq\widetilde{\sigma}>0,$ and $\nu\in[4,\infty]$, we have $\sup_{f\in\mathcal{F}}P|f|^{k}\leq\widetilde{\sigma}^{2}b^{k-2}$ for $k=2,3$ and $\|F\|_{P,\nu}\leq b$. Then for every $\gamma\in(0,1)$ there exists a random variable $\mathbf{S}\overset{d}{=}\sup_{f\in\mathcal{F}}G_{P}f$ such that
$\mathbb{P}\Big(\Big|\sup_{f\in\mathcal{F}}\mathbb{G}_{n}f-\mathbf{S}\Big|>\frac{b\Gamma_{n}}{\gamma^{1/2}n^{1/2-1/\nu}}+\frac{(b\widetilde{\sigma})^{1/2}\Gamma_{n}^{3/4}}{\gamma^{1/2}n^{1/4}}+\frac{(b\widetilde{\sigma}^{2}\Gamma_{n}^{2})^{1/3}}{\gamma^{1/3}n^{1/6}}\Big)\leq C\Big(\gamma+\frac{\log n}{n}\Big)$
where $)\Gamma_{n}=cv(\log n\vee\log(Ab/\widetilde{\sigma}$)), and $c,C>0$ are constants that depend only on $\nu$ ($1/\nu$ is interpreted as 0 when $\nu=\infty$).
Acknowledgments. The authors are grateful to the Editor-in-Chief, Dietrich von Rosen, an Associate Editor and the reviewers for their valuable comments, which led to an improved manuscript. Moreover, we thank Max Farrell for his insightful remarks.